OpenStudy (anitas.):

Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = −2 x = 0 x = 2 There is no x value, as there is no solution. x can be any value, as there is an infinite number of solutions.

Directrix (directrix):

1) 2x + y = 1 2) 4x + 2y = −2

Directrix (directrix):

In equation #1, solve for y. 2x + y = 1 -2x -2x ------------- y = 1 - 2x

Directrix (directrix):

In equation #2, replace the variable y by what you got when you solved for y in equation #1. 4x + 2y = −2 4x + 2*(1 - 2x) = -2

Directrix (directrix):

Crank that out and solve for x. @AloneS

OpenStudy (anitas.):

ghh im not alones

Directrix (directrix):

Pardon me. @AnitaS. Crank that out and solve for x. 4x + 2*(1 - 2x) = -2

OpenStudy (anitas.):

ok I get 2=-2

Directrix (directrix):

I got 0 = -4. There is no solution for this pair of equations. Hold on while I draw the graphs. I think that the lines are parallel.

OpenStudy (anitas.):

i got no solution too

OpenStudy (photon336):

yeah when you solve a system of equations and you end up with different number's on opposite sides of the equal sine that means no solution exists. for example 2 = 3

OpenStudy (photon336):

sign***

Directrix (directrix):

Because the lines are parallel, they do not intersect and there is no common solution. @AnitaS.

Directrix (directrix):

That is a wrap.

OpenStudy (anitas.):

So the answer is C right?

OpenStudy (anitas.):

ty! ;)

Directrix (directrix):

If this is C, then yes. There is no x value, as there is no solution.